OpenStudy (anonymous):

What is the area of a circle with a radius of 5 meters? Use twenty-two sevenths for pi. nine and twenty-three twenty-eighths square meters nineteen and nine-fourteenths square meters thirty-nine and two-sevenths square meters seventy-eight and four-sevenths square meters

OpenStudy (anonymous):

Quite a few variables.... I really don't feel like doing this for you. Equation: A=πr^2 π=Θ

OpenStudy (anonymous):

Wait.... Whoops read it wrong... Solve this: 22/7(5)^2

OpenStudy (anonymous):

seventy-eight and four-sevenths square meters

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

22/7 * 5^2 = 550/7 = 78 + 4/7

OpenStudy (anonymous):

Isn't that exactly what I said xD ^ <3